aA + bB -> cC

whose first and second half-lives are 49.6 and 49.6 minutes respectively. If the rate constant is equal to 0.01397 and [A]0 = 10.05, calculate the [A] at t = 21.78 minutes.
Answer #1


In[A] - ln [A] = kt in 10.05 - Im CA] = 0.01397% 21.78 in CA] : Im 1005 -0.01391 % 21.78 In (A) = 2.307 - 0-304 In TA] = 2.003 TA] 282003 CA] = "7. 411 2 LAJ = Concentration OdA at t= 21 78 minutes